OpenStudy (anonymous):

A sporting goods store ordered an equal number of pairs of ski boots for males and females. the factory delivered 50 extra pairs of female boots by mistake, making the ratio of male boots to female boots 5 to 7. how many pairs of ski boots did the store originally order

OpenStudy (anonymous):

250

OpenStudy (anonymous):

work please?

OpenStudy (anonymous):

well 5 to 5 is basically 1 to 1 so when the ratio then became 5 to 7 with fifty extra means that for every 1 increase in the ratio an extra 25 was added. or 5/7 = x/(x+50) 5x + 250 = 7x 2x = 250 x = 125 so 125 per gender of boot